OpenStudy (anonymous):

in a conference each journalist shakes each journalist's hand. it was total 182 shakes. how many journalist were in that conference?

OpenStudy (anonymous):

and please give an explanation

OpenStudy (across):

\[handshakes=\frac{1}{2}(n)(n-1)\]where n is the number of journalists. Rearrange the equation and find your solution.

OpenStudy (anonymous):

no... the answer is 19 and your given equation doesnt work

OpenStudy (across):

...

OpenStudy (across):

\[182=\frac{1}{2}(n)(n-1)\]

OpenStudy (anonymous):

yes i understand that, but the answer is n=19, but it doesnt work from your equation! 19*18

OpenStudy (anonymous):

19*18*0.5=171

OpenStudy (across):

20*19*0.5=190

OpenStudy (anonymous):

so?

OpenStudy (across):

Are you sure there were 182 handshakes?

OpenStudy (anonymous):

yes, definately.. and i am sure that the answer is 19

OpenStudy (whpalmer4):

So, do you disagree with the formula, or with the evaluation of 1/2*19*18 = 171? The solution of 1/2*(n-1)*n = 182 is n = 1/2(1+sqrt(1457)) journalists...

OpenStudy (whpalmer4):

if you have A, B, and C, A shakes B, A shakes C, B shakes C are the 3 unique handshakes. For A, B, C, D it is A shakes B, C and D, B shakes C, D, C shakes D for 6 unique handshakes. For A, B, C, D, E it is A shakes B,C,D,E, B shakes C, D, E, C shakes D, E, D shakes E for 10 unique handshakes. the number of handshakes is always of the form 1 + 2 + 3 + 4 + ... and there is no such sum that equals 182. sum from 1 to 18 is 171, sum from 1 to 19 is 190.

OpenStudy (anonymous):

well, i think the same, that's why I wanted to ask if I am right... maybe there is a mistake in my book...

OpenStudy (across):

Perhaps your book is being "realistic" and knows that not everyone in the room could possibly shake hands. If that's the case, then we need at least 20 people in the room as 19 people will uniquely shake hands up to 171 times (assuming someone didn't shake hands more than once).

OpenStudy (across):

Other than that, your book has an error.

OpenStudy (anonymous):

hahahaha you are right thanks anyway!